OpenStudy (anonymous):

i need help with this long devision Algebra b problem (a squared plus 6a) divided by (a+6)

OpenStudy (anonymous):

\[\frac{a^2+6a}{a+6}=\frac{a\left(a+6\right)}{a+6}=a\]Voila!

OpenStudy (anonymous):